and collected A research company conducted a domogranhic information on the respondents from each country Some Primary Post-A research company conducted a survey and collected demographic information on the respondents from each country. To the right is a table showing the number of people with various levels of education in five countries. Complete parts​ a) through​ d).
Answer #1

a)P(from country E) =1552/7677 =0.2022

b)P(completed his or her education before college )=(4193+1177)/7677 =0.6995

c)P(from B or post graduate) =(1538+356-65)/7677 =0.2382

d)

P(B and primary school or less)=(316/7677)=0.0412